# Numerical computation of Stephenson’s $g$-functions in multiply connected domains

### v1

@@ 1 Introduction
For $E_{r}\neq\emptyset$ and $\partial\Omega_{r}\backslash E_{r}\neq\emptyset$, the harmonic measure of $E_{r}$ with respect to $\Omega_{r}$ is the $C^{2}(\Omega_{r})$ function $u:\Omega_{r}\to(0,1)$ satisfying the Laplace equation
@@
in $\Omega_{r}$, with $u(z)\to 1$ when $z\to E_{r}$ and $u(z)\to 0$ when $z\to\partial\Omega_{r}\backslash E_{r}$. Harmonic measure is a key concept in potential theory and has numerous applications to geometric function theory [1, 4, 6, 11, 27]. The harmonic measure of $E_{r}$ with respect to $\Omega_{r}$ calculated at the point $z\in\Omega_{r}$ will be denoted by $\omega_{\rm hm}(z,E_{r},\Omega_{r})$.
@@
where $g(0)=0$. The $g$-function, which was first introduced in [3, Problem 6.116], is a non-decreasing piecewise continuous function. It follows from this definition of the $g$-function that $g(r)=u(z_{0})$ where $u(z)$ is the unique solution of the following Dirichlet boundary value problem (BVP):
@@
$\gamma(z)=\left\{\begin{array}[]{cc}1,&z\in E_{r},\\ 0,&z\in\partial\Omega_{r}\backslash E_{r}.\\ \end{array}\right.$ (3)
@@
The $h$-function can be computed by solving a Dirichlet boundary value problem similar to (2). This is illustrated also in Figure 1 for unbounded multiply connected domains $\Omega$ and in Figure 2 for bounded domains.
@@
Despite this body of work on the $h$-functions, the study of $g$-functions has so far received much less attention. Analytic formulas for the $g$-function for several simply connected domains has been presented recently in [16, 19]. In this paper, we present a boundary integral equation method for the numerical computation of the $g$-function in multiply connected domains. The proposed method is used effectively in this paper to make calculations of the $g$-functions associated with multiply connected circular and rectilinear slit domains. To the best of our knowledge, this is the first attempt to compute the $g$-function for multiply connected domains.
@@ 2 The integral equation method
where $\Gamma_{0}$ is the external boundary and oriented counterclockwise. The inner curves $\Gamma_{1},\ldots,\Gamma_{m}$ are oriented clockwise. Each curve $\Gamma_{k}$ is parametrized by $\eta_{k}(t)$ for $t\in J_{k}=[0,2\pi]$, $k=0,1,\ldots,m$. If $\Gamma_{k}$ has corner points (but not cusps), we parametrize it as explained in [24]. Let $J$ be the disjoint union of the $m+1$ intervals $J_{j}=[0,2\pi]$, $j=0,1,\ldots,m$. We define a parametrization of the whole boundary $\Gamma$ on $J$ by
$\eta(t)=\left\{\begin{array}[]{cc}\eta_{0}(t),&t\in J_{0},\\ \eta_{1}(t),&t\in J_{1},\\ \vdots\\ \eta_{m}(t),&t\in J_{m}.\\ \end{array}\right.$
@@
For each $j$, $j=0,1,\ldots,m$, there exists a unique real function $\mu_{j}(t)$ and a unique piecewise constant function $\nu_{j}(t)$ [25, Theorem 2]:
@@
$\gamma(t)=\left\{\begin{array}[]{cc}0,&t\in J_{0},\\ 1,&t\in J_{1},\\ \vdots\\ 1,&t\in J_{m}.\\ \end{array}\right.$ (16)
@@
where $\gamma_{0}(t)=\gamma(t)$. Here, the right-hand side is a piecewise constant function which implies that $\gamma_{0}(t)+\nu_{0}(t)=0$ and hence the right-hand side of the linear system (15) will be the vector $[0,1,\ldots,1]^{T}$.
@@ 3.2.3 $\Omega_{r}$ is multiply connected
For $11\leq m\leq\ell-$ 1 and $r\in(|z_{m+1}|-R_{m+1},|z_{m+1}|+R_{m+1})$, the capture circle $\mathcal{C}_{r}$ intersects the circle $C_{m+1}$ at the two points $\xi_{1}$ and $\xi_{2}$ given by (18). Note that the circles $C_{1},\ldots,C_{m}$ are inside $\mathcal{C}_{r}$ and the circles $C_{m+2},\ldots,C_{\ell}$ are outside $\mathcal{C}_{r}$.
@@ 4.1 Continuous boundary data
$\hat{\gamma}(z)=\left\{\begin{array}[]{cc}0,&z\in\Gamma_{0},\\ 1,&z\in\Gamma_{j},\quad j=1,\ldots,m.\\ \end{array}\right.$ (28)
@@ 4.2.1 The function $V_{r}$
$\gamma(z)=\left\{\begin{array}[]{cc}1,&z\in I_{k}^{\prime},\\ 0,&z\in\mathcal{C}_{r}.\\ \end{array}\right.$ (30)
@@
which transplants $\hat{\Omega}_{r}$ onto the upper half-plane such that $I_{k}^{\prime}$ is mapped onto the finite slit $[-w_{k},w_{k}]$ on the real line and $\mathcal{C}_{r}$ is mapped onto ({})\$({\mathbb{R}}\cup\{\infty\})\backslash[-w_{k},w_{k}]$, where
@@ 4.2.3 $\Omega_{r}$ is multiply connected
For $11\leq m\leq\ell-$ 1 and $r\in(a_{m+1},b_{m+1})$, the capture circle $\mathcal{C}_{r}$ intersects the slit $I_{m+1}$ at the point $c_{m+1}=r$. Note that the slits $I_{1},\ldots,I_{m}$ are inside $\mathcal{C}_{r}$ and the slits $I_{m+2},\ldots,I_{\ell}$ are outside $\mathcal{C}_{r}$. In this case, the domain $\Omega_{r}$ is a bounded multiply connected domain of connectivity $m+1$ and
@@
$\hat{\gamma}(z)=\left\{\begin{array}[]{cc}1-V_{r}(z),&z\in I_{1}\cup\cdots\cup I_{m},\\ 0,&z\in I_{m+1}^{\prime}\cup{\mathcal{C}}_{r}.\\ \end{array}\right.$ (36)
@@
$\hat{\gamma}(\zeta)=\left\{\begin{array}[]{cc}1-V_{r}(\psi^{-1}(\Phi(\zeta))),&\zeta\in\Gamma_{1}\cup\cdots\cup\Gamma_{m},\\ 0,&\zeta\in\Gamma_{0}.\\ \end{array}\right.$ (39)
@@ 4.3 Numerical examples
We assume that the basepoint is $z_{0}=0$ and $\Omega_{k}$ is the unbounded multiply connected domain in the exterior of the closed set $E_{k}$ for $k=0,1,2,3,4$ where $E_{0}=[1,8]$, $E_{1}=[1,2]$, $E_{2}=[1,2]\cup[3,4]$, $E_{3}=[1,2]\cup[3,4]\cup[5,6]$, and $E_{4}=[1,2]\cup[3,4]\cup[5,6]\cup[7,8]$. The values of the $g$-function $g(r)$, for $k=0,1,2,3,4$, computed with $n=2^{13}$ are shown in Figure 8.
Figure 8: The values of the $g$-function $g(r)$ for $\Omega_{k}$ in Example 1 for $k=0,1,2,3,4$.
@@
is the middle-thirds Cantor set of the interval $[1,2]$. We assume that the basepoint is $z_{0}=0$ and $\Omega_{k}$ is the unbounded multiply connected domain in the exterior of the closed set $E_{k}$. The graphs of the $g$-function $g(r)$, for $k=0,2,4,8,16,32$, computed with $n=2^{13}$ are shown in Figure 9. These graphs illustrate that the $g$-function for the domain $\Omega_{k}$ is non-constant and exhibits a finite number of points where its first derivative is discontinuous. Note that, for this domain $\Omega_{k}$, the graphs of the $h$-function for $k=16$ and $k=32$ are presented in [8, Fig 8]. In contrast, these graphs consist of a finite number of horizontal line segments which correspond to those values of $r$ when the capture circle is not intersecting the slits and is otherwise an increasing function.
@@ 5 Other domains
[…]
@@
$U(w)=\left\{\begin{array}[]{cc}1,&\zeta\in I_{1}\cup\cdots\cup I_{\ell},\\ 0,&\zeta\in\hat{I}_{1}\cup\cdots\cup\hat{I}_{\ell}.\\ \end{array}\right.$
@@ 6 Bounded domains
To complete our study of the $g$-function in this paper, it is worth demonstrating that the presented method can be used in a straightforward manner for bounded domains $\Omega$. This is briefly described in this section through an example for the domains $\Omega$ in Figure 6 but make them bounded domains with adding an outer boundary component which is the circle with center $z_{0}=0$ and radius $R_{0}=12$. For $120\leq r<$ 12, the domain $\Omega_{r}$ here is the same as in the unbounded domain case in Figure 6. Hence, the values of $g(r)$ for $120\leq r<$ 12 are the same as the values for the unbounded case as presented in Figure 7. For all values of $r$ such that $r\geq 12$, we will have $\Omega_{r}=\Omega$ and hence $g(r)=1$. The graphs of the $g$-function $g(r)$ for the three cases with $\ell=10$ are shown in Figure 11. All graphs exhibit a fixed jump when $r=12$.
Figure 11: The values of the $g$-function $g(r)$ for the three cases in Figure 6 made bounded domains by adding an outer boundary component $|z|=12$.
@@ 7 Conclusion
This work is the first attempt at computing the $g$-function in multiply connected domains. The graphs of the $g$-functions presented in this paper provide some insight into the properties of these $g$-functions and how they compare to the graphs of their $h$-function counterparts shown in [7, 8, 9]. However, much numerical work still remains to be done; in particular, to compute $g$-functions associated with domains bounded by curves other than circles and rectilinear slits, and for different basepoint locations.

### v2

@@ 1 Introduction
[…]
@@
in $\Omega_{r}$, with $u(z)\to 1$ when $z\to E_{r}$ and $u(z)\to 0$ when $z\to\partial\Omega_{r}\backslash E_{r}$. Harmonic measure is a key concept in potential theory and has numerous applications to geometric function theory [1, 4, 5, 6, 11, 27]. The harmonic measure of $E_{r}$ with respect to $\Omega_{r}$ calculated at the point $z\in\Omega_{r}$ will be denoted by $\omega_{\rm hm}(z,E_{r},\Omega_{r})$.
@@
where $g(0)=0$. The $g$-function, which was introduced by Stephenson in [3, Problem 6.116], is a non-decreasing piecewise continuous function. It follows from this definition of the $g$-function that $g(r)=u(z_{0})$ where $u(z)$ is the unique solution of the following Dirichlet boundary value problem (BVP):
@@
[…]
@@
The $h$-function can be computed by solving a Dirichlet boundary value problem similar to (2). This is illustrated also in Figure 1 for unbounded multiply connected domains $\Omega$ and in Figure 2 for bounded domains. The $h$-function was also introduced by Stephenson in [3] and was first studied in depth by Walden & Ward [28].
@@
Despite this body of work on the $h$-functions, the study of $g$-functions has so far received much less attention. Analytic formulas for the $g$-function for several simply connected domains have been presented recently in [16, 19]. In this paper, we present a boundary integral equation method for the numerical computation of the $g$-function in multiply connected domains. The proposed method is used effectively in this paper to make calculations of the $g$-functions associated with multiply connected circular and rectilinear slit domains. To the best of our knowledge, this is the first attempt to compute the $g$-function for multiply connected domains.
@@ 2 The integral equation method
where $\Gamma_{0}$ is the external boundary and oriented counterclockwise. The inner curves $\Gamma_{1},\ldots,\Gamma_{m}$ are oriented clockwise. Each curve $\Gamma_{k}$ is parametrized by $\eta_{k}(t)$ for $t\in J_{k}=[0,2\pi]$, $k=0,1,\ldots,m$. We assume that $\eta_{k}(t)$ is twice continuously differentiable such that $\eta^{\prime}_{k}(t)\neq 0$ for $t\in J_{k}$. If $\Gamma_{k}$ has corner points (but not cusps), we parametrize it as explained in [24]. Let $J$ be the disjoint union of the $m+1$ intervals $J_{j}=[0,2\pi]$, $j=0,1,\ldots,m$. We define a parametrization of the whole boundary $\Gamma$ on $J$ by
[…]
@@
For $j=0,1,\ldots,m$, there exists a unique real function $\mu_{j}(t)$ and a unique piecewise constant function $\nu_{j}(t)$ [25, Theorem 2]:
@@
[…]
@@
where $\gamma_{0}(t)=\gamma(t)$. Here, the right-hand side is a piecewise constant function which implies that $\gamma_{0}(t)+\nu_{0}(t)=0$ and hence the right-hand side of the linear system (15) will be the vector $0,1,\ldots,1]^{T}$.
@@ 3.2.3 $\Omega_{r}$ is multiply connected
For $1\leq m\leq\ell-1$ and $r\in(|z_{m+1}|-R_{m+1},|z_{m+1}|+R_{m+1})$, the capture circle $\mathcal{C}_{r}$ intersects the circle $C_{m+1}$ at the two points $\xi_{1}$ and $\xi_{2}$ given by (18). Note that the circles $C_{1},\ldots,C_{m}$ are inside $\mathcal{C}_{r}$ and the circles $C_{m+2},\ldots,C_{\ell}$ are outside $\mathcal{C}_{r}$.
@@ 4.1 Continuous boundary data
[…]
@@ 4.2.1 The function $V_{r}$
[…]
@@
which transplants $\hat{\Omega}_{r}$ onto the upper half-plane such that $I_{k}^{\prime}$ is mapped onto the finite slit $w_{k},w_{k}]$ on the real line and $\mathcal{C}_{r}$ is mapped onto $({\mathbb{R}}\cup\{\infty\})\backslash[-w_{k},w_{k}]$, where
@@ 4.2.3 $\Omega_{r}$ is multiply connected
For $1\leq m\leq\ell-1$ and $r\in(a_{m+1},b_{m+1})$, the capture circle $\mathcal{C}_{r}$ intersects the slit $I_{m+1}$ at the point $c_{m+1}=r$. Note that the slits $I_{1},\ldots,I_{m}$ are inside $\mathcal{C}_{r}$ and the slits $I_{m+2},\ldots,I_{\ell}$ are outside $\mathcal{C}_{r}$. In this case, the domain $\Omega_{r}$ is a bounded multiply connected domain of connectivity $m+1$ and
@@
[…]
@@
[…]
@@ 4.3 Numerical examples
We assume that the basepoint is $z_{0}=0$ and $\Omega_{k}$ is the unbounded multiply connected domain in the exterior of the closed set $E_{k}$ for $k=0,1,2,3,4$ where $E_{0}=[1,8]$, $E_{1}=[1,2]$, $E_{2}=[1,2]\cup[3,4]$, $E_{3}=[1,2]\cup[3,4]\cup[5,6]$, and $E_{4}=[1,2]\cup[3,4]\cup[5,6]\cup[7,8]$. The values of the $g$-function $g(r)$, for $k=0,1,2,3,4$, computed with $n=2^{13}$, are shown in Figure 8 (left). A magnification of the part of the figure on $2,6]\times[0.3,0.5]$ is shown in Figure 8 (right).
The sets $E_{0}$ and $E_{1}$ consist of only one interval and the graphs of the $g$-function $g(r)$ for these two sets only have one corner point at which $g^{\prime}(r)$ is discontinuous. This occurs at $r=1$ when the capture circle first hits each of the sets $E_{0}$ and $E_{1}$. The set $E_{2}$ consists of two intervals and the graph of the $g$-function $g(r)$ for this set shows that $g^{\prime}(r)$ is discontinuous at both $r=1$ and $r=3$ (i.e. at the left-ends of the two intervals making up the set $E_{2}$). Similarly, the graphs of the $g$-function $g(r)$ for the sets $E_{3}$ and $E_{4}$ show that $g^{\prime}(r)$ is discontinuous at $r=1,3,5$ for $E_{3}$ and at $r=1,3,5,7$ for $E_{4}$.
Refer to caption Refer to caption
Figure 8: The values of the $g$-function $g(r)$ for $\Omega_{k}$ in Example 1 for $k=0,1,2,3,4$ (left) and the magnification of the part of the figure on $2,6]\times[0.3,0.5]$ (right).
@@
is the middle-thirds Cantor set of the interval $[1,2]$. We assume that the basepoint is $z_{0}=0$ and $\Omega_{k}$ is the unbounded multiply connected domain in the exterior of the closed set $E_{k}$. The graphs of the $g$-function $g(r)$, for $k=0,2,4,8,16,32$, computed with $n=2^{13}$ are shown in Figure 9. These graphs illustrate that the $g$-function for the domain $\Omega_{k}$ is non-constant and exhibits a finite number of points where its first derivative is discontinuous. These points correspond to the values of $r$ at the $2^{k}$ left-ends of the intervals making up the set $E_{k}$. Note that, for this domain $\Omega_{k}$, the graphs of the $h$-function for $k=16$ and $k=32$ are presented in [8, Fig 8]. In contrast, these graphs consist of a finite number of horizontal line segments which correspond to those values of $r$ when the capture circle is not intersecting the slits and is otherwise an increasing function.
@@ 5 Other domains
where $\zeta^{\ast}=\zeta_{\lfloor(3\ell/2)+1\rfloor}$ and $\lfloor\cdot\rfloor$ denotes the floor function. The point $\zeta^{*}$ is chosen such that it will be mapped onto $1$ on the positive real line so that the arc connecting $\zeta_{1}$ and $\zeta_{\ell+1}$ (in the counterclockwise direction) is mapped onto the negative real line. Then, for $j=2,\ldots,2\ell$, the points $x_{j}=\phi(\zeta_{j})$ will be on the real line such that $x_{j}<x_{j+1}$. Let $I_{j}$ and $\hat{I}_{j}$ be the images of the arcs $L_{j}$ and $\hat{L}_{j}$ under the Möbius transformation $w=\phi(\zeta)$, respectively. Then $I_{2},\ldots,I_{\ell},\hat{I}_{1},\ldots\hat{I}_{\ell-1}$ are finite intervals and $I_{1},\hat{I}_{\ell}$ are infinite intervals on the real line.
@@
[…]
@@ 6 Bounded domains
To complete our study of the $g$-function in this paper, it is worth demonstrating that the presented method can be used in a straightforward manner for bounded domains $\Omega$. This is briefly described in this section through the following example. Let $\Omega$ be the bounded domain lying in the interior of the circle with center $z_{0}=0$ and radius $R_{0}=12$ and exterior to the $\ell$ circles with the centers $z_{k}$ and radii $R_{k}$, $k=1,\ldots,\ell$, for the three cases considered in Section 3.3. Then $\Omega$ is a bounded multiply connected domain of connectivity $11$. For $0\leq r<12$, the domain $\Omega_{r}$ here is the same as in the unbounded domain case in Figure 6. Hence, the values of $g(r)$ for $0\leq r<12$ are the same as the values for the unbounded case as presented in Figure 7. For all values of $r$ such that $r\geq 12$, we will have $\Omega_{r}=\Omega$ and hence $g(r)=1$. The graphs of the $g$-function $g(r)$ for the three cases with $\ell=10$ are shown in Figure 11. All graphs exhibit a fixed jump when $r=12$.
Figure 11: The values of the $g$-function $g(r)$ for the bounded domain $\Omega$ when $\ell=10$.
@@ 7 Conclusion
This work is the first attempt at computing the $g$-function in multiply connected domains. The graphs of the $g$-functions presented in this paper provide some insight into the properties of these $g$-functions and how they compare to the graphs of their $h$-function counterparts shown in [7, 8, 9]. However, much numerical work still remains to be done; in particular, to compute $g$-functions associated with domains bounded by curves other than circles and rectilinear slits, and for different basepoint locations. Further, Stephenson’s original questions in [3] are mainly concerned about the connection between the geometry of the domain $\Omega$ and the behavior of both the $g$-function and the $h$-function. These questions are still not fully answered and hence provide a source of future research topics.
Acknowledgments
The authors would like to thank an anonymous reviewer for their valuable comments and suggestions. CCG acknowledges the support of a LEAPS-MPS grant from the National Science Foundation (#2532158).